Brest State Technical University

(http://www.bstu.unibel.by/index-e.html)

Administration

• Rector Petr C. Poita (phone:42-7457)

• Vice-Rector Vyacheslav I. Dragan (phone:42-0361)

• Vice-Rector Timofei N Bazenkov (phone:42-2052)

• Vice-Rector Viktor B. Tur (phone:42-4271)

• Vice-Rector Slavii I. Efimenko (phone:42-1852)